OpenStudy (anonymous):

A researcher studying the nutritional value of a new candy places a 3.00-gram sample of the candy inside a bomb calorimeter and combusts it in excess oxygen. The observed temperature increase is 2.09 °C. If the heat capacity of the calorimeter is 46.30 kJ·K–1, how many nutritional Calories are there per gram of the candy?

OpenStudy (zale101):

Change the temperature from celsius to kelvins so you can use the specific heat capacity equation \(\Large Q=c*\Delta T\) So, \(\Large \Delta T_{celsius}=\Delta T_{kelvins}\)

OpenStudy (zale101):

because this involves specific heat capacity

OpenStudy (zale101):

When calculating Q=m*c*deltaT, the only thing that matters with temperature is the difference or temperature change. The size of the temperature change is the same as the temperature change in kelvins.

OpenStudy (zale101):

2.09 °C this is just a temperature change in celsius so, changed into kelvins 2.09 k

OpenStudy (zale101):

@carlos1291996

OpenStudy (anonymous):

yes

OpenStudy (zale101):

\(Q = c\Delta T \) c is kj delta t is the delta K Q will be in kJ plug everything and tell me what you get for kJ

OpenStudy (zale101):

dont forget, nutritional Calories are in kcal. So after you get kJ, you convert it to kcal

OpenStudy (anonymous):

So it would be (46.30X10^3)/(2.09) then devide that by 3?

OpenStudy (anonymous):

i am sorry i am really confuse :/

OpenStudy (zale101):

Q=c * delta T Q= (46.30 kJ K^-1)*(2.09 K) Q=96.77 kJ Then convert from kj to kcal. How many kJ are there for kcal?

OpenStudy (zale101):

Then reason i said you should convert kj to kcal, because the question is asking "how many \(\color{red}{nutritional~Calories}\) are there per gram of the candy?" The question wants the answer in kcal per grams. Normally, the calories we see on food labels (aka. nutritional calories) is usually in 1000 calories, which is kilocalories (kcal).

OpenStudy (zale101):

Here's a hint: \(\LARGE 1~~kcal=4.184 ~~kJ\)

OpenStudy (zale101):

Use this as a conversion factor to convert from kJ to kcal

OpenStudy (zale101):

makes sense? @carlos1291996

OpenStudy (anonymous):

so 96.77/4.184=23.12kcal

OpenStudy (zale101):

Yep :)

OpenStudy (zale101):

We are not done yet, we only got the first part of the answer, getting kcal's but the question wants nutritional calories per \(\color{red}{grams}\)

OpenStudy (zale101):

How are you going to get the kcal into kcal per grams (aka. \(\LARGE \frac{kcal}{grams}\))

OpenStudy (zale101):

@carlos1291996

OpenStudy (anonymous):

I got it ! it's (46.30 kj*C) X ( 2.09 C ) / 3.00 g X 0.239 kcal/kj

OpenStudy (anonymous):

just want to make sure is it correct?

OpenStudy (anonymous):

The answer i got was 7.71 Cal/g

OpenStudy (zale101):

Yes, to get kcal per grams You have 23.15 for kcal and 3.00 for grams \(\LARGE \frac{kcal}{grams}=\frac{23.15~kcal}{3.00~grams}=\large 7.71~ kcal ~~per ~~grams\)

OpenStudy (zale101):

Good job ^_^ @carlos1291996

OpenStudy (zale101):

typo, i meant 23.12 not 23.15

OpenStudy (anonymous):

oh okay thank you!

OpenStudy (zale101):

You're welcome! :)